### v2

# cadrille: Multi-modal CAD Reconstruction with Online Reinforcement Learning

Abstract
[…]
Computer-Aided Design (CAD) plays a central role in engineering and manufacturing, making it possible to create precise and editable 3D models. Using a variety of sensor or user-provided data as inputs for CAD reconstruction can democratize access to design applications. However, existing methods typically focus on a single input modality, such as point clouds, images, or text, which limits their generalizability and robustness. Leveraging recent advances in vision-language models (VLM), we propose a multi-modal CAD reconstruction model that simultaneously processes all three input modalities. Inspired by large language model (LLM) training paradigms, we adopt a two-stage pipeline: supervised fine-tuning (SFT) on large-scale procedurally generated data, followed by reinforcement learning (RL) fine-tuning using online feedback, obtained programatically. Furthermore, we are the first to explore RL fine-tuning of LLMs for CAD tasks demonstrating that online RL algorithms such as Group Relative Preference Optimization (GRPO) outperform offline alternatives. In the DeepCAD benchmark, our SFT model outperforms existing single-modal approaches in all three input modalities simultaneously. More importantly, after RL fine-tuning, cadrille sets new state-of-the-art on three challenging datasets, including a real-world one.
1 Introduction
Computer-Aided Design (CAD) is the core of modern engineering and manufacturing, providing the tools to create detailed and modifiable 3D models briere2012comparing-3d-cad . Creating CAD models manually requires skills, time, and effort. To simplify this process, CAD reconstruction aims at generating CAD models directly from scanned objects, making the process faster, cheaper, and more accessible overall rukhovich2024cad-recode .
Typically, CAD models are created with a sequence of 2D sketches and 3D operations willis2021fusion360 ; wu2021deepcad . This representation allows CAD models to be easily edited, making it prevalent in popular CAD tools like SolidWorks and AutoCAD and in CAD generation research. Most existing CAD generation methods define CAD sequences using special command tokens khan2024cad-signet ; wu2021deepcad . However, state-of-the-art results are obtained via mapping CAD sequences to casual Python code rukhovich2024cad-recode . Following the same paradigm, we generate CAD models as executable Python scripts.
The most well-studied input modality in CAD reconstruction is naturally a point cloud rukhovich2024cad-recode . However, point clouds can only be obtained when a physical 3D object is available, while scanning usually requires special equipment, making the process complicated for non-experts. Images capture finer details and can be sourced using customer low-end devices (e.g., smartphone cameras), hence relaxing the hardware requirements chen2025cadcrafter ; chen2024img2cad . In the meantime, textual descriptions can enrich the object representation with semantic context khan2024text2cad . Using various input modalities, such as multi-view images, or natural language descriptions, would make design assistance applications simple even for non-experienced users. However, existing approaches typically focus on a single modality at a time, limiting their robustness and generality.
Figure 1: Compared to CAD-Recode, the only existing method of CAD reconstruction as LLM-based code generation, cadrille has two key novelties. First, it goes beyond the standard training scheme and pioneers LLM RL fine-tuning for CAD reconstruction (left). Moreover, besides point clouds only accepted by single-modal CAD-Recode, cadrille extends to images and textual descriptions, making it the first multi-modal approach delivering state-of-art results (right).
The task of Python code generation given various inputs naturally refers to the large Vision-Language models (VLM), which demonstrate strong reasoning capabilities across diverse modalities shao2024deepseekmath ; wang2024qwen2-vl ; yang2024qwen2 . In this work, we harness the power of VLMs to build a multimodal CAD reconstruction model that simultaneously processes point clouds, multi-view images, and texts, and generates Python code.
Existing CAD reconstruction methods face generalization issues due to how they are trained. Specifically, handcrafted CAD datasets are small and limited in diversity, while models trained with procedurally generated data struggle to transfer to the real-world domain khan2024cad-signet ; wu2021deepcad . Inspired by the standard LLM training pipelines shao2024deepseekmath , we introduce a multi-stage training paradigm in the context of CAD reconstruction. We use voluminous procedurally generated data for supervised training, while valuable but scarcer handcrafted data is reserved for RL fine-tuning. This scheme eliminates the need for large-scale handcrafted data and allows the model to first generalize across the CAD domain and then specialize using preference-based objectives. We test Direct Preference Optimization (DPO) rafailov2023dpo , which was employed for the CAD reconstruction in  chen2025cadcrafter . Moreover, since online RL techniques dominate in tasks where feedback can be programatically computed, we also try the online Group Relative Preference Optimization (GRPO) technique. Our experiments demonstrate that online RL is more effective for CAD reconstruction, which aligns with the results obtained in other domains shao2024deepseekmath .
Our experiments show that cadrille  outperforms existing modality-specific baselines in accuracy. Moreover, RL fine-tuning ensures validity of generated Python code, which posed a challenge for prior works. As a result, the proposed approach demonstrates unattainable robustness and sets a new state-of-the-art on several CAD datasets, including a real-world CC3D mallis2023cc3d . Essentially, this opens up new possibilities for generalization in open-world scenarios.
In summary, our contributions are as follows:
  • We propose cadrille, a CAD reconstruction model that operates jointly on point cloud, images, and text modalities;
  • We are the first to explore RL fine-tuning of LLM models in the context of CAD reconstruction. Specifically, we propose a novel paradigm to split procedurally generated and handcrafted data between supervised fine-tuning (SFT) and RL fine-tuning stages;
  • With a single model, we simultaneously achieve state-of-the-art reconstruction quality for all input modalities on the DeepCAD dataset. RL fine-tuning pushes the quality even further in three benchmarks, including real-world CC3D.
2 Related Work
CAD Generation
Existing CAD generation methods can be classified into three categories based on CAD model representations: constructive solid geometry (CSG)  du2018inversecsg ; ellis2019write ; friedrich2019optimizing ; kania2020ucsg-net ; nandi2018functional ; ren2021csg-stump ; sharma2018csgnet ; tian2019learning ; yu2023d2csg ; yu2022capri-net , boundary representation (B-rep) guo2022complexgen ; jayaraman2022solidgen ; lambourne2021brepnet ; li2019supervised ; li2025caddreamer ; liu2024nvdnet ; liu2024point2cad ; sharma2020parsenet ; smirnov2021learning ; wang2022neural ; wang2020pie-net ; xu2024brepgen and CAD sequence badagabettu2024query2cad ; chen2024img2cad ; chen2025cadcrafter ; dupont2024transcad ; khan2024cad-signet ; khan2024text2cad ; lambourne2022prismcad ; ma2024cad-diffuser ; mallis2024cad-assistant ; ren2022extrudenet ; rukhovich2024cad-recode ; wang2025cad-gpt ; wu2021deepcad ; xu2022skexgen ; xu2023hnc-cad ; yuan2025cad-editor ; zhang2024flexcad . In the CSG foley1996csg paradigm, CAD is represented as a CSG tree constructed using boolean operations (union, subtraction, difference) of geometric primitives (e.g., cubes, cylinders, or spheres). This approach fails to express intricate shapes and is generally not well-aligned with how engineers and designers actually build CAD models. B-rep ansaldi1985brep is a graph that describes connections between faces, edges, and vertices of a 3D model. Creating a B-Rep requires enforcing topological consistency on edges, which introduces additional complexity to the generation procedure and complicates editing of generated models.
CAD Sequence Reconstruction
Unlike general CAD generation, which may prioritize plausibility, diversity, or creativity in design, CAD reconstruction aims at faithfulness to the given inputs, requiring the output model to match the original shape.
Point clouds are the most well-studied input modality in CAD reconstruction. The seminal work on point cloud-based CAD reconstruction by Wu et al. wu2021deepcad proposed encoding CAD sketch-and-extrude sequences as special tokens. Beyond that, DeepCAD, a large-scale dataset of 180k hand-crafted CAD models, was presented. Subsequent works dupont2024transcad ; khan2024cad-signet ; xu2023hnc-cad adopted the same CAD representation and trained on the same DeepCAD dataset. More recently, CAD-Recode rukhovich2024cad-recode introduced a paradigm shift by representing CAD models as Python code, providing greater expressiveness and flexibility, and released a new training dataset of approximately 1 million procedurally generated CAD samples.
Only few recent works chen2025cadcrafter ; chen2024img2cad ; khan2024text2cad ; wang2025cad2program have explored CAD reconstruction from alternative input modalities, such as single- or multi-view images and natural language descriptions. These approaches extend the DeepCAD dataset by rendering synthetic views or generating textual captions for existing CAD models. Among them, CADCrafter chen2025cadcrafter stands out for its unified framework that handles both single- and multi-view inputs, whether rendered or real. In the domain of text-to-CAD, the most notable results come from Text2CAD khan2024text2cad , which uses a vision-language model (VLM) to generate detailed captions for DeepCAD shapes and then trains a model to predict the corresponding sketch-and-extrude sequences from these descriptions.
Generally, state-of-the-art CAD reconstruction approaches are tailored to process specific input modalities with distinct architectures, while multimodal CAD reconstruction remains underexplored. Recent CAD-GPT wang2025cad-gpt reconstructs a CAD model given a single image and textual description, yet falls behind CADCrafter chen2025cadcrafter and Text2CAD khan2024text2cad by a large margin (up to two orders of magnitude!). We claim cadrille to be the first competitive multimodal CAD reconstruction approach, which can jointly handle three input modalities (point cloud, images, and text) within a unified framework.
RL for CAD and LLM
The training of modern LLMs wang2024qwen2-vl ; yang2024qwen2 is performed in several stages. The process starts with unsupervised pretraining on internet-scale data to gain broad structural knowledge of the textual domain. Then, supervised fine-tuning (SFT) is applied to train the model for specific tasks. Alignment, also known as preference optimization, helps to strengthen model’s focus on the best answers, which can be sourced from human feedback, or automatically by using elaborate RL approaches ouyang2022rlhf ; rafailov2023dpo ; shao2024deepseekmath .
Reinforcement learning has also been explored in CAD-related tasks chen2025cadcrafter ; sharma2018csgnet ; yin2025rlcad ; chao2025reinforcement . For example, RLCAD yin2025rlcad proposes an RL-based policy that interacts with a CAD engine to generate sketch-and-revolve sequences, receiving feedback based on the quality of reconstruction. Chao et. al. chao2025reinforcement applies Deep Q-Networks to recover parametric CAD models from 2D orthographic drawings in the form of loop-paths. CADCrafter chen2025cadcrafter implements a code checker to serve as an implicit reward model and fine-tunes using offline Direct Preference Optimization (DPO) rafailov2023dpo .
However, none of these works applies reinforcement learning to fine-tune large language models (LLMs) for CAD reconstruction. In this work, we are the first to explore RL fine-tuning of LLMs in this domain. We show that online RL methods like GRPO shao2024deepseekmath are particularly effective, eliminating the need for human feedback.
3 CAD Sequence Reconstruction
Problem Formulation
The task of CAD reconstruction implies recovering a CAD model given a multimodal input $q$, which can be a 3D point cloud, a set of images, or a textual description. We represent CAD models as Python scripts rukhovich2024cad-recode that, when executed, generate a parametric Boundary Representation (B-Rep) of a 3D shape. Respectively, given an input $q$, we search for a trainable policy $\pi_{\theta}$, s. t. $\pi_{\theta}(q)$ produces a token sequence $\tau$, which is essentially a text of a Python program generating a CAD model.
Multimodal Data
Figure 2: Overview of multimodal data generation pipeline producing textual descriptions, multi-view images and point clouds.
For training a model, we derive all input modalities from ground-truth CAD models (Fig. 2). Below, we describe how each modality is constructed according to the established data generation protocols chen2025cadcrafter ; khan2024text2cad ; wu2021deepcad .
Given a CAD model as a parametric 3D shape (B-Rep), we sample points directly from the parametric surfaces of the model. Modern CAD engines provide built-in routines for surface sampling, making it simple and straightforward.
To generate images, the B-Rep is first tessellated, i.e., converted into a triangle mesh that approximates the surface geometry. Then, this mesh can be rendered from multiple viewpoints to produce multi-view image inputs.
Generating textual data is notably more challenging. Since our goal is accurate geometry reconstruction rather than generating a semantically relevant sample, inputs should provide detailed and comprehensive geometric information. Consequently, loose textual descriptions are generally insufficient. The necessary level of granularity is investigated in Text2CAD khan2024text2cad , where LLMs and VLMs are combined in a multi-stage sophisticated pipeline that generates textual descriptions from both the CAD sequence and rendered images.
4 Proposed Method
4.1 cadrille Architecture
Figure 3: Overview of cadrille. It can handle three input modalities within a unified framework. Point clouds are processed with a trainable projection layer, while images and texts are passed to a VLM directly. The output of the model is an executable Python script for CAD generation.
The architecture of our cadrille is depicted in Fig. 3. The model accepts inputs in the form of a point cloud, a set of images, or a text prompt, and outputs a Python code, that when executed, produces a CAD model. cadrille is build on top of a VLM that provides native support of text and image inputs and is already capable of generating Python code. Textual input is passed through the original embedding layer, and images are processed with an original visual encoder. The point cloud processing logic is the same as in CAD-Recode rukhovich2024cad-recode . Specifically, we use a single projection layer to embed 3D points, sample points from the surface via furthest point sampling, and do not use normals.
4.2 Supervised Fine-tuning
As shown on Fig. 1, cadrille benefits from three stages of training. First, we use VLM which is pre-trained on the internet-scale data in the unsupervised manner. After this stage, VLM is able to process textual and visual inputs and generate Python code, but lacks mechanisms to handle point clouds. In this work, we do not perform any unsupervised VLM training, but enjoy the capabilities of an already trained model.
The second stage is supervised fine-tuning for a specific task. During SFT, a model develops the ability to process point clouds and learns a policy $\pi_{\theta}$ to map multimodal inputs $q$ to Python codes $\tau$, making SFT an essential part of cadrille pipeline. We construct a training dataset $\mathcal{D}$ of samples $(q,\ \tau)$, where $q$ is a multimodal input. The training procedure aims to minimize cross-entropy between ground truth and predicted Python code tokens:
$\mathbb{E}_{(q,\tau)\sim\mathcal{D}}\left[\log\pi_{\theta}(\tau\mid q)\right]$
4.3 Limitations of SFT
Two-stage training has already been adopted in CAD-Recode rukhovich2024cad-recode , which employs supervised fine-tuning (SFT) to adapt a pretrained language model for point cloud-based CAD reconstruction. However, this strategy reveals its limitations in a cross-domain scenario: CC3D IoU is as low as 60% and the invalidity ratio (IR) is to 10%, which means that every tenth prediction fails to produce a valid output (Tab. 3, row 2). To mitigate this issue, CAD-Recode uses a test-time sampling technique. For each input query, 10 candidate Python programs are generated, and the candidate with the highest IoU is selected. After that, IoU increases to 74%, while IR drops below 0.5%. However, this improvement comes at the cost of a 10x increase in inference time. Can similar gains be achieved without sacrificing test-time efficiency?
To maintain fast and simple inference, we shift our focus to improving the training process. Training solely on procedurally generated CAD data might limit performance in real-world applications. Nevertheless, training on handcrafted models also presents challenges, e.g., prior work rukhovich2024cad-recode shows that SFT directly on the DeepCAD dataset harms performance, leading to a 10% drop in IoU.
Our experiments confirm that simply mixing procedurally generated and handcrafted data for training fails to improve results and can even degrade performance (Tab. 3, row 4). We attribute this to inconsistency in CAD sequences across datasets: for instance, DeepCAD models are constructed using commands like extruded cuts and symmetric extrusions, which are not present in the generation procedure of the CAD-Recode dataset.
To address this limitation, we introduce a novel third stage in the training pipeline, namely, reinforcement learning fine-tuning on handcrafted data not annotated with CAD sequences. This approach resolves inconsistency issues while still allowing the model to adapt to real-world domain.
4.4 RL Fine-tuning
We formulate RL fine-tuning as follows. Given a dataset of inputs (either images or point clouds) $\mathcal{D}=\{q_{i}\}_{i=1}^{N}$, and reward function $R(\tau)$, we learn LLM policy $\pi_{\theta}(\tau\mid q)$ that generates a Python code $\tau$ for an input $q$, s.t. it maximizes the expected reward $\mathbb{E}_{q_{i}\sim D,\tau_{i}\sim\pi_{\theta}(\cdot\mid q_{i})}\left[R(\tau_{i})\right]$.
Note that at this stage annotated pairs of $(q,\tau)$ are not needed for supervision, since Python codes $\tau$ are being sampled from the trained SFT model. In fact, CAD sequences are not needed for RL fine-tuning, and the data requirements can be relaxed to 3D meshes instead. This is especially beneficial from the practical perspective, as RL fine-tuning can be performed using generally more accessible mesh datasets, which opens new possibilities for training models accommodated to artifacts present in real-world data.
The reward function $R(\tau)$ is a combination of terms that address precision and robustness:
$R(\tau)=r_{\text{IoU}}(\tau)+r_{\text{invalid}}(\tau),$
where $r_{\text{IoU}}$ is an IoU between the CAD model produced by $\tau$ and ground truth 3D mesh, additionally multiplied by a factor of 10 to enforce precise reconstruction. $r_{\text{invalid}}$ penalizes invalid predictions: it is set to -10 for invalid $\tau$ and 0 otherwise.
Empirically, we found that hard example mining leads to a faster convergence of RL fine-tuning. Consequently, we only use examples $q$ where the reward $R(\tau)$ averaged over three samples produced by the SFT model is less than 7.5.
DPO
Direct Preference Optimization (DPO) rafailov2023dpo learns from pairwise preference data, approximating an implicit reward via a reparameterized Bradley-Terry model.
We construct the training dataset by sampling $K=5$ Python codes $\tau$ for each input $q$ from the SFT model $\pi_{\theta_{r}}$ at temperature $T=1.0$. At each training step for the given sample, we randomly select two outputs. The output with a larger reward $R(\tau)$ is considered to be a preferred prediction $\tau_{w}$, and another is non-preferred $\tau_{l}$. The optimization objective is formulated as:
$\mathbb{E}_{(q,\tau_{w},\tau_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta_{t}}(\tau_{w}\mid q)}{\pi_{\theta_{r}}(\tau_{w}\mid q)}-\beta\log\frac{\pi_{\theta_{t}}(\tau_{l}\mid q)}{\pi_{\theta_{r}}(\tau_{l}\mid q)}\right)\right]$
DPO training starts with $\pi_{\theta_{r}}$ and proceeds for 10 epochs. After that, the SFT model is replaced with the latest $\pi_{\theta_{t}}$, and trained for another 10 epochs. In this way, the model gradually diverges from the original SFT model. In our experiments, we found it to be beneficial for performance.
However, DPO performance is upper-bounded by the quality of the best generated sample for a given example. This limitation cannot be overcome without generating additional samples, so we adapt an online RL approach that can benefit from newly generated samples.
Dr. CPPO
We combine two recent modifications of the popular GRPO shao2024deepseekmath method: Dr. GRPO liu2025dr.grpo which eliminates the need for a reference model and modifies the objective, and CPPO lin2025cppo which uses samples with the strongest signal. The hybrid approach ensures both computational efficiency and accuracy; hereinafter, it is referred to as Dr. CPPO.
$G$ sequences $\{\tau_{g}\}_{g=1}^{G}$ are sampled from the current policy $\pi_{\theta_{\text{old}}}(\tau\mid q)$ for a given input $q$ with temperature $T=1.0$. For each output $g$, the advantage $A_{g}$ is estimated as $A_{g}={r_{g}-\operatorname{mean}(\{r_{i}\}_{i=1}^{G})}$. $N$ samples with the highest $|A_{g}|$ are used to form a batch $\mathcal{B}$ and perform policy update by maximizing PPO schulman2017ppo objective:
$\mathbb{E}_{\{\tau_{g}\}\sim\mathcal{B}}\left[\min\left(\frac{\pi_{\theta_{t}}(\tau_{g}\mid q)}{\pi_{\theta_{\text{old}}}(\tau_{g}\mid q)}A_{g},\ \text{clip}\left(\frac{\pi_{\theta_{t}}(\tau_{g}\mid q)}{\pi_{\theta_{\text{old}}}(\tau_{g}\mid q)},\ 1-\epsilon,\ 1+\epsilon\right)A_{g}\right)\right]$
5 Experiments
Datasets
DeepCAD wu2021deepcad (denoted as D in Tables) serves as our primary benchmark for supervised training. We adopt the Text2CAD version of DeepCAD, which enriches it with textual descriptions. The training set comprises approximately 160k samples, while 8046 are left for testing.
For SFT, we also use the procedurally generated CAD-Recode rukhovich2024cad-recode dataset (denoted as R). It is an order of magnitude larger than DeepCAD, consisting of approximately 1 million CAD programs written in CadQuery cadquery , a parametric Python-based CAD language.
Fusion360 willis2021fusion360 (denoted as F) is a small CAD reconstruction benchmark with complex and realistic CAD models. In the standard experimental protocol, only the test subset (1725 samples) is used, as the lack of Python CAD sequences makes it unsuitable for conventional supervised training. Still, we can use its training set (6900 samples) for our annotation-free RL fine-tuning.
To show the versatility and applicability of our approach, in addition to handcrafted and procedurally generated meshes, we report metrics on the CC3D dataset mallis2023cc3d . It features 2973 input point clouds sampled from real scans of CAD models with noisy values, missing parts, and smoothed edges.
Metrics
Following CAD-Recode rukhovich2024cad-recode , we evaluate the quality of the predicted CAD models using three metrics: Chamfer Distance (CD), Intersection over Union (IoU), and Invalidity Ratio (IR). Since invalid CAD models introduce a notable bias into mean estimates, we report both mean and a more robust median CD, both computed using 8192 points. CD values are multiplied by 103. The divergence between ground truth and reconstructed meshes is measured using IoU (in %). The IR indicates the percentage of generated sequences that do not produce a valid CAD model.
5.1 Supervised Fine-Tuning
Results on DeepCAD
In Tab. 1, we compare cadrille with single-modality CAD reconstruction methods on DeepCAD. Here, we specify input modalities with subscripts: p stands for point clouds, i denotes images, and t is for texts. cadrille trained jointly on point clouds, multi-view images, and texts from the DeepCAD training set (Dpit) outperforms the modality-specific baselines. Noticeably, IR is reduced almost twice for point clouds (from 1.1 to 0.4) and 7x for images (3.6 to 0.5).
Training using the large-scale procedurally generated CAD-Recode dataset (R) consistently improves accuracy over training on the DeepCAD dataset. Since we also use a Qwen LLM model as in CAD-Recode (Qwen2-VL-2B against Qwen2-1.5B), comparable quality of point cloud-based reconstruction is expected. When training cadrille on point clouds and images (Rpi), it maintains the same accuracy on point clouds but additionally extends to images. After training with point clouds, images, and texts (Spi + Di), cadrille generalizes across modalities without loss of quality on each modality. For fair comparison, we do not apply any RL techniques in this series of experiments, and mix up training datasets trivially for SFT.
Results on Fusion360 and CC3D
Both Fusion360 and CC3D datasets do not provide annotations in a compatible format, and are only used for testing in the standard evaluation protocol khan2024cad-signet . Accordingly, testing on these datasets is performed in a zero-shot scenario, which allows assessing the generalization ability of CAD reconstruction approaches. Furthermore, since CC3D contains real scans of objects, this experiment emulates real-world application.
We report quality of image-based and point cloud-based CAD reconstruction in Tab. 2 and  3, respectively. CADCrafter chen2025cadcrafter is the only method performing CAD reconstruction based on multi-view images. However, the authors of CADCrafter only report metrics on the DeepCAD dataset, and benchmarking it on other datasets is problematic since the code has never been released. To establish a baseline in image-based CAD reconstruction, we combine two off-the-shelf state-of-the-art methods, namely, multi-view reconstruction method LRM hong2024lrm ; xu2024instantmesh and CAD-Recode. LRM takes multi-view images as inputs and produces a mesh, which is turned into a point cloud via surface sampling, and this point cloud is then passed to CAD-Recode to create a CAD model. As can be seen in Tab. 2, cadrille  trained on CAD-Recode outperforms both baselines.
In Tab. 3, we compare cadrille  against the state-of-the-art approaches originally trained on the DeepCAD (CAD-SIGNet khan2024cad-signet ) and CAD-Recode rukhovich2024cad-recode datasets. As could be expected, cadrille  is on par with CAD-Recode, while delivering substantially better quality w.r.t. CAD-SIGNet.
Method Train Point Cloud Multi-view Images Text
Data CD↓ IoU↑ IR↓ CD↓ IoU↑ IR↓ CD↓ IoU↑ IR↓
PointNet→DeepCAD wu2021deepcad Dp 9.64 46.7 7.1
Point-BERT→HNC-CAD xu2023hnc-cad Dp 8.64 65.3 5.6
MultiCAD ma2023multicad Dp 8.09 11.5
TransCAD dupont2024transcad Dp 4.51 65.5 1.1
PrismCAD lambourne2022prismcad Dp 4.28 72.1 16.2
Point2Cyl uy2022point2cyl Dp 4.27 73.8 3.9
CAD-Diffuser ma2024cad-diffuser Dp 3.02 74.3 1.5
CAD-SIGNet khan2024cad-signet Dp 0.29 77.3 5.0
DINOv2→HNC-CAD xu2023hnc-cad Di 2.08 10.1
DINOv2→DeepCAD wu2021deepcad Di 1.13 10.6
CADCrafter chen2025cadcrafter Di 0.26 3.6
BERT→DeepCAD wu2021deepcad Dt 32.82 10.0
Text2CAD khan2024text2cad Dt 0.37 71.5 0.9
cadrille Dpit 0.25 79.4 0.4 0.25 78.2 0.5 0.21 81.1 1.4
CAD-Recode rukhovich2024cad-recode Rp 0.18 87.1 3.1
cadrille Rpi 0.18 87.1 2.1 0.18 86.1 1.5
cadrille Rpi+Dt 0.18 87.1 2.1 0.18 86.1 1.5 0.20 82.1 1.4
Table 1: Results on DeepCAD test set. The best results are bold, the second best are underlined. Our cadrille trained jointly on three modalities outperforms all existing modality-specific methods. Here, we report metrics obtained without RL fine-tuning or test-time sampling for fair comparison.
Method RL Train Data DeepCAD Fusion360 CC3D
SFT RL CD↓ IoU↑ IR↓ CD↓ IoU↑ IR↓ CD↓ IoU↑ IR↓
LRM→CAD-Recode Rp 0.53 69.8 14.3 0.62 62.5 18.7 1.19 50.1 20.1
CADCrafter chen2025cadcrafter DPO Di Di 0.26 3.6
cadrille Rpi 0.18 86.1 1.5 0.20 77.6 3.2 0.81 56.1 7.7
cadrille Rpi+Dpi 0.19 85.6 0.6 0.23 75.2 2.6 1.17 53.1 6.0
cadrille DPO Rpi $\text{D}^{\text{--}}_{\text{i}}\text{+}\text{F}^{\text{--}}_{\text{i}}$ 0.18 86.9 1.8 0.20 78.5 1.7 0.89 56.0 3.9
cadrille Dr. CPPO Rpi $\text{D}^{\text{--}}_{\text{i}}\text{+}\text{F}^{\text{--}}_{\text{i}}$ 0.17 92.2 0.0 0.17 84.6 0.0 0.57 65.0 0.1
Table 2: Results of CAD reconstruction from multi-view images. With RL fine-tuning, cadrille achieves best results across three benchmarks.
Method RL Train Data DeepCAD Fusion360 Real-world CC3D
SFT RL CD↓ IoU↑ IR↓ CD↓ IoU↑ IR↓ CD↓ IoU↑ IR↓
CAD-SIGNet khan2024cad-signet Dp 0.29 77.3 5.0 0.70 58.4 9.3 4.42 39.1 15.5
CAD-Recode rukhovich2024cad-recode Rp 0.18 87.1 3.1 0.19 79.1 5.0 0.54 60.5 9.8
cadrille Rpi 0.18 87.1 2.1 0.19 79.8 2.8 0.54 61.8 5.9
cadrille Rpi+Dpi 0.19 86.6 0.9 0.22 76.5 2.0 0.79 58.7 4.1
cadrille DPO Rpi $\text{D}^{\text{--}}_{\text{i}}\text{+}\text{F}^{\text{--}}_{\text{i}}$ 0.18 88.1 0.7 0.19 80.9 1.3 0.54 61.3 2.6
cadrille Dr. CPPO Rpi $\text{D}^{\text{--}}_{\text{i}}\text{+}\text{F}^{\text{--}}_{\text{i}}$ 0.17 90.2 0.0 0.17 85.0 0.2 0.47 67.9 0.2
Table 3: Results of CAD reconstruction from point clouds. cadrille performs on par with CAD-Recode rukhovich2024cad-recode when trained on the CAD-Recode dataset (R). With RL, cadrille establishes state-of-the-art on DeepCAD, Fusion360 and real-world CC3D.
DeepCAD Dataset Fusion360 Dataset Real-world CC3D Dataset
Point Cloud Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
CAD-SIGNet Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
CAD-Recode Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
cadrille Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: CAD models reconstructed from point clouds from the DeepCAD, Fusion360, and CC3D datasets.
DeepCAD Dataset Fusion360 Dataset CC3D Dataset
Images Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LRM→ CAD-Recode Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
cadrille Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: CAD models reconstructed from multi-view images on the DeepCAD, Fusion360, and CC3D datasets.
5.2 Reinforcement Learning
RL on single modality boosts other modalities
In Tab. 2, we report accuracy of image-based CAD reconstruction on three benchmarks. Respectively, we fine-tune cadrille  with images from DeepCAD and Fusion360 datasets. It is worth noticing that while Fusion360 cannot be used for direct supervised training, it can still contribute to RL fine-tuning where CAD sequences are not required, so we can benefit from adding it to the mixture. We denote DeepCAD and Fusion360 datasets without CAD sequence annotations as $\text{D}^{\text{--}}$ and $\text{F}^{\text{--}}$.
Surprisingly, RL fine-tuning on images appears to be beneficial for other modalities: as reported in Tab. 3, the model tuned on $\text{D}^{\text{--}}_{\text{i}}\text{+}\text{F}^{\text{--}}_{\text{i}}$ (row 6) delivers state-of-the-art quality of CAD reconstruction from point clouds as well.
RL improves metrics in cross-dataset scenario
RL fine-tuning with DeepCAD and Fusion360 boosts accuracy on the test splits of the respective datasets. Yet, the performance gain is not limited to the domains seen by a model during SFT and RL fine-tuning. In image-based CAD reconstruction, CD is reduced from 0.81 to 0.57, while IR dropped dramatically from 7.7 to 0.1 (rows 3 and 6, respectively). When testing on point clouds, RL also improves all scores on CC3D, making IR less than 0.2%, which is negligible.
Online RL outperforms offline RL
Fine-tuning cadrille  using offline DPO reduces IR twice in most cases, while accuracy scores are not affected (rows 3 and 5 in both Tables). In the meantime, Dr. CPPO beats SFT in terms of all metrics, adding 3-9% to IoU scores and bringing IR under 0.2% in all benchmarks (row 6). The observed improvement of CAD reconstruction accuracy aligns well with the experimental results obtained in other tasks where feedback can be programmatically computed shao2024deepseekmath .
RL fine-tuning beats SFT on a mixture
A common assumption is that mixing datasets improves generalization by increasing data diversity and volume. However, our experiments show that SFT with a plain mixture of CAD-Recode and DeepCAD datasets (Rpi+Dpi, row 4) does not lead to performance gains, and can even degrade results w.r.t. SFT with Rpi (row 3). We attribute this effect to the domain gap between datasets, specifically, some CAD operations present in DeepCAD (e.g., symmetric extrusion, extruded cut) are lacking from CAD-Recode.
Qualitative results
CAD models obtained with RL fine-tuning are depicted in Fig. 4 (from point clouds) and Fig. 5 (from multi-view images). Compared to predecessors, cadrille  produces more geometrically plausible reconstructions and better restores fine details.
6 Limitations
Despite strong performance, our method has several limitations. It relies on either relatively small and limited hand-crafted benchmarks (DeepCAD), or large-scale procedurally generated datasets (CAD-Recode), which may not reflect the complexity of real-world CAD data. Due to the unavailability of CC3D’s training split in the public domain, training or annotation-free fine-tuning on real data remains underexplored. Additionally, the text modality is underutilized, since natural language descriptions are scarce, and VLM-generated captions from Text2CAD often lack realism and might impose risks of data leakage due to being overly descriptive and detailed. Moreover, while cadrille can process all three modalities, it treats them independently and lacks mechanisms to compensate for low-quality or missing inputs. Besides, we only use images for RL fine-tuning. Tuning on point clouds appeared to be unstable, which we attribute to 1) point clouds being a non-native input modality for an LLM and 2) suboptimal point cloud processing strategies inherited from CAD-Recode.
7 Conclusion
We introduced cadrille, a multimodal CAD reconstruction model that is capable of processing point clouds, multi-view images, and text inputs within a unified VLM-based framework. By adopting a two-stage training paradigm, namely, supervised fine-tuning on synthetic data followed by reinforcement learning fine-tuning with programmatic feedback, we improved both reconstruction quality and validity ratio. Our empirical study demonstrated that online RL approaches are especially beneficial in the CAD reconstruction scenario. The proposed method achieves new state-of-the-art results in multiple benchmarks, including a real-world dataset, highlighting its robustness, generalizability, and potential for further use in applications.
Appendix A Qualitative Results
Real-world single-image experiment
Since our cadrille  is trained only on synthetic renders, sim-to-real transfer can be eligibly questioned. To address the potential concerns about the applicability of our approach, apart from validating on real scans from the CC3D dataset, we also experiment with CAD reconstruction from a single real-world image.
The pipeline consists of three steps. First, an input image is processed using the recent image-to-mesh InstantMesh xu2024instantmesh method, that produces a mesh. Second, we follow the same protocol as in other experiments to convert this mesh to a point cloud 6 or four multi-view images 7. We claim that the obtained results look promising, and this practical pipeline opens new opportunities of in-the-wild CAD reconstruction.
Input image InstantMesh (mesh) Point cloud cadrille (CAD)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 6: Results of CAD reconstruction from a single real-world image. cadrille  takes point clouds sampled from mesh reconstructed by InstantMesh as input.
Input image InstantMesh (mesh) 4-view images cadrille (CAD)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 7: Results of CAD reconstruction from a single real-world image. cadrille  takes multi-view images rendered from mesh reconstructed by InstantMesh as input.
Text-based CAD reconstruction
Fig. 8 shows results of text-based CAD reconstruction on the DeepCAD dataset. Long textual descriptions still cannot define even simple 3D shapes comprehensively and unambiguously, making CAD reconstruction from textual inputs the most challenging. Both Text2CAD khan2024text2cad and cadrille struggle to recover correct geometry, yet our approach yields more accurate predictions, which is also reflected in the quantitative metrics.
Text Text2CAD cadrille GT
Create a new coordinate system with Euler angles set to 0 degrees for the first 2 angles and -90 degrees for the third angle. Set the translation vector to 0.2766, 0.1476, and 0.2766. On the first face, draw a 2-dimensional sketch consisting of 2 loops. For the first loop, draw a circle with its center at 0.0984, 0.0984, and a radius of 0.0984. For the second loop, draw another circle with the same center but a smaller radius of 0.048. Scale the entire 2-dimensional sketch by a factor of 0.1969. Rotate and translate the scaled 2-dimensional sketch using the previously defined coordinate system settings. Extrude the outer loop (the larger circle) by 0.1476 units in the direction opposite to the normal. Similarly, extrude the inner loop (the smaller circle) by the same distance in the same direction. Ensure that the extrusion results in a new solid body. The final dimensions of the cylindrical object with a hole should be a length of 0.1969 units, a width of 0.1969 units, and a height of 0.1476 units. The object resembles a doughnut shape, with a smooth surface and a uniform diameter. Refer to caption Refer to caption Refer to caption
Create a new coordinate system with Euler Angles set to [0, 0, 0] and a Translation Vector of [0, 0, 0]. Draw a 2D sketch on the XY plane and define the first face using a loop of 4 lines. The first line starts at (0, 0) and ends at (0.5, 0). The second line starts at (0.5, 0) and ends at (0.5, 0.75). The third line starts at (0.5, 0.75) and ends at (0, 0.75). The fourth line starts at (0, 0.75) and ends at (0, 0). Scale the 2D sketch using a scale factor of 0.75. Transform the scaled 2D sketch into 3D using the same Euler Angles and Translation Vector. Extrude the 2D sketch to create a 3D model with an extrusion depth of 0.5 units towards the normal. Create a new solid body with this extrusion and verify the dimensions: length of 0.5 units, width of 0.75 units, and height of 0.5 units. Next, create a new coordinate system with Euler Angles set to [0, 0, 0] and a Translation Vector of [0.1, 0, 0.5]. Draw a 2D sketch on the XY plane and define the first face using a loop of 4 lines. The first line starts at (0, 0) and ends at (0.3, 0). The second line starts at (0.3, 0) and ends at (0.3, 0.75). The third line starts at (0.3, 0.75) and ends at (0, 0.75). The fourth line starts at (0, 0.75) and ends at (0, 0). Scale the 2D sketch using a scale factor of 0.75. Transform the scaled 2D sketch into 3D using the same Euler Angles and Translation Vector. Extrude the 2D sketch to create a 3D model with an extrusion depth of 0.4 units in the opposite direction of the normal. Remove material from the existing body using this extrusion and verify the dimensions: length of 0.3 units, width of 0.75 units, and height of 0.4 units. The final shape is a U-shaped bracket with a rectangular cross-section. It has 2 parallel sides and an open space in the middle. The dimensions are: length of 0.5 units, width of 0.75 units, and height of 0.5 units. Refer to caption Refer to caption Refer to caption
Create a new coordinate system with Euler angles set to 0, 0, and -90 degrees, and a translation vector of 0, 0.2812, and 0.1406. On the first face, draw the first loop as a rectangle with the following lines: the first line starts at (0, 0) and ends at (0.3398, 0); the second line starts at (0.3398, 0) and ends at (0.3398, 0.4687); the third line starts at (0.3398, 0.4687) and ends at (0, 0.4687); the fourth line starts at (0, 0.4687) and ends at (0, 0). In the second loop, draw a circle with a center at (0.2344, 0.2344) and a radius of 0.0937. In the third loop, draw a circle with a center at (0.2344, 0.082) and a radius of 0.0176. In the fourth loop, draw a circle with a center at (0.2344, 0.3867) and a radius of 0.0176. Apply a scale factor of 0.4687 to the entire sketch. Rotate the scaled sketch using the Euler angles set in the coordinate system and translate it using the translation vector. Extrude the sketch 0.2812 units along the normal direction without extruding in the opposite direction to create a new solid body. The final dimensions of the rectangular box with circular cutouts are a length of 0.3398 units, a width of 0.4687 units, and a height of 0.2812 units. Refer to caption Refer to caption Refer to caption
Figure 8: Results of text-based CAD reconstruction on the DeepCAD dataset.
Input Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
cadrille Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Failure cases of CAD reconstruction from point clouds and multi-view images on DeepCAD, Fusion360, and CC3D datasets.
Failure cases
are depicted in Fig. 9. cadrille  always predicts a geometrically relevant shape, but still might miss details, especially for objects with complex and granular surfaces.
Appendix B Quantitative Results
Data for RL fine-tuning
As described in Sec. 4.4 in the main paper, not all available data is used for RL fine-tuning, but only the most hard examples are considered. By varying the threshold, we can adjust the difficulty level and the ratio of data selected for fine-tuning. Finding the proper balance is crucial for RL fine-tuning, since it has a notably larger time- and memory footprint compared to SFT, and is hardly feasible without hard example mining.
RL Data DeepCAD Fusion360 CC3D
DeepCAD Fusion360 CD↓ IoU↑ IR↓ CD↓ IoU↑ IR↓ CD↓ IoU↑ IR↓
10k 0.7k 0.18 89.0 1.0 0.18 81.0 2.6 0.74 59.8 6.3
20k 1.4k 0.17 90.4 0.7 0.18 82.3 0.9 0.63 62.4 4.0
30k 2k 0.17 90.9 0.2 0.17 83.5 0.5 0.62 62.3 0.9
40k 2.5k 0.17 91.2 0.1 0.17 83.9 0.2 0.60 63.3 0.7
50k 3k 0.17 92.2 0.0 0.17 84.6 0.0 0.57 65.0 0.1
Table 4: Results of CAD reconstruction from multi-view images with varying amount of data used for RL fine-tuning.
In Tab. 4, we report results obtained when fine-tuned on data of different volume. All metrics improve gradually with an increase of the amount of the training data. Still, all results reported in this table supersede the SFT baseline (Tab. 2 of the main paper).
Mean CD
Mean CD is sometimes reported alongside median CD. In the main paper, we only provided median CD as the most common CAD reconstruction metric. Nevertheless, the evaluation protocol might be considered incomplete without mean CD. In Tab. 5, we report mean CD metric for all scenarios on all three datasets. Our SFT model outperforms all competitors. The gain is especially tangible in text-based CAD reconstruction, where Text2CAD mean CD is reduced 6x from 26.4 to 3.95. With RL fine-tuning (Dr. CPPO) the new state-of-the-art is set in both image-based and point cloud-based CAD reconstruction on all three datasets. The most dramatic improvement is demonstrated on the Fusion360 dataset, where the relative increase exceeds 40% (Tab. 6).
Method Train RL DeepCAD Fusion360 CC3D
Data Points Images Text Points Images Points Images
PointNet→DeepCAD wu2021deepcad Dp 42.5 330
CAD-SIGNet khan2024cad-signet Dp 6.81 14.5 32.6
CAD-Recode rukhovich2024cad-recode Rp 0.83 1.21 3.21
LRM→CAD-Recode Rp 3.36 4.33 4.75
BERT→DeepCAD wu2021deepcad Dt 97.9
Text2CAD khan2024text2cad Dt 26.4
cadrille Dpit 3.43 3.57 4.24 7.61 8.59 12.2 13.2
cadrille Rpi+Dt 0.76 0.81 3.95 1.10 1.13 2.32 3.50
cadrille Rpi DPO 0.61 1.35 0.84 1.27 2.32 3.33
cadrille Rpi Dr. CPPO 0.57 0.43 0.58 0.64 1.86 2.68
Table 5: Mean CD scores obtained across all benchmarks and available input modalities. RL fine-tuning is performed using $\text{D}^{\text{--}}_{\text{i}}\text{+}\text{F}^{\text{--}}_{\text{i}}$ data.
More baselines on Fusion360
In Tab. 6, we compare cadrille  to more point cloud-based CAD reconstruction baselines on the Fusion360 dataset. As can be observed, cadrille  outperforms competitors trained either on DeepCAD or CAD-Recode datasets.
Inference time
When target at practical use, efficiency is an issue. In Tab. 7, we compare inference time of our multi-modal cadrille  against previous best single-modal methods: CAD-Recode rukhovich2024cad-recode , Text2CAD khan2024text2cad , and our baseline LRM→CAD-Recode.
Our cadrille  is built on top of Qwen2-VL-2B, the smallest Qwen2 model with vision capabilities. When inferred on point clouds, it is 20% slower compared to CAD-Recode using Qwen2-1.5B. The image inference takes comparable time to proceed. Processing text prompts from Text2CAD dataset lasts notably longer, so that the inference time almost doubles and reaches 3.9 seconds. Text2CAD uses a smaller and faster BERT-large model, that allows achieving efficiency at cost of accuracy. Compared to Text2CAD, cadrille is delivers 6x better mean CD 5 being only 2x slower.
Method Train CD↓ IoU↑ IR↓
DeepCAD wu2021deepcad Dp 89.2 39.9 25.2
MultiCAD ma2023multicad Dp 42.2 16.5
HNC-CAD xu2023hnc-cad Dp 36.8 63.5 7.3
TransCAD dupont2024transcad Dp 33.4 60.2 2.4
CAD-Diffuser ma2024cad-diffuser Dp 3.85 63.2 1.7
CAD-SIGNet khan2024cad-signet Dp 0.70 58.3 9.3
cadrille Dp 0.66 63.7 0.6
CAD-Recode rukhovich2024cad-recode Rp 0.19 79.1 5.0
cadrille Rp 0.19 79.8 2.8
Table 6: Results of point-based CAD reconstruction on the Fusion360 test set. All reported metrics are obtained using an SFT model without RL.
Method LLM Points Images Text
CAD-Recode rukhovich2024cad-recode Qwen2-1.5B 1.8
LRM→CAD-Recode Qwen2-1.5B 4.4
Text2CAD khan2024text2cad BERT-336M 1.7
cadrille Qwen2-VL-2B 2.0 2.0 3.9
Table 7: Inference time in seconds measured on the DeepCAD dataset. All methods are benchmarked on the single H100 GPU with a batch size of 1.
Single-image CAD reconstruction
To compare against single-image CAD reconstruction methods, namely, CAD-GPT wang2025cad-gpt and Img2CAD chen2024img2cad , we conduct an experiment with single-view images on the DeepCAD test set. Our SFT model improves over CADCrafter chen2025cadcrafter , reducing median CD from 0.72 to 0.21. As could be expected, the results of single-view CAD reconstruction are slightly inferior to the ones obtained from multi-view images (81% vs 86% IoU).
Method CD↓ IoU↑ IR↓
GPT-4o hurst2024gpt-4o ; wang2025cad-gpt 62.6 64.4
CAD-GPT wang2025cad-gpt 9.77 1.6
DINOv2→HNC-CAD xu2023hnc-cad 2.14 11.4
Img2CAD chen2024img2cad 1.60 28.8
DINOv2→DeepCAD wu2021deepcad 1.26 12.3
CADCrafter chen2025cadcrafter 0.72 8.1
cadrille 0.21 81.7 1.3
Table 8: Results of CAD reconstruction from a single image on the DeepCAD dataset. All reported metrics are obtained with an SFT model without RL.
Method CD↓ IoU↑ IR↓
GPT-o4-mini 2.37 60.4 15.9
cadrille 0.17 92.2 0.0
Table 9: Results of CAD reconstruction from multi-view images on the DeepCAD dataset.
Zero-shot CAD reconstruction
Image-based CAD reconstruction can be performed in zero-shot way using existing VLMs. CAD-GPT wang2025cad-gpt sets a weak baseline using GPT-4o, that has an invalidity ratio of 64% Tab. 8
We construct another baseline for CAD reconstruction from multi-view images. Four images rendered with orthogonal viewing directions are given to GPT-o4-mini to produce a Python code of CAD model. We apply iterative closest point (ICP) to align predictions before computing metrics so that correct predictions with wrong orientation are not penalized. As can be seen in Tab.  9, this strategy allows achieving significantly better results compared to CAD-GPT. Still, invalidity ratio is as high as 15% and IoU is 30% lower compared to our cadrille.
Appendix C Implementation Details
Architecture
Our model is built on top of Qwen2-VL wang2024qwen2-vl and uses its native capabilities of image and text understanding. In all experiments on multi-view image CAD reconstruction, we use four images. Images are rendered with fixed camera positions, and concatenated into 2x2 grid, forming a combined image of size 268 $\times$ 268 px (as shown in Fig. 7 and  9). This combined image is passed through the Qwen vision encoder, that outputs 400 input tokens. The point cloud injecting into LLM is implemented exactly as in CAD-Recode rukhovich2024cad-recode . Specifically, our input consists of 256 unordered 3D points without normals, sampled from the surface using the furthest point sampling method. The points are projected into shared embedding space with a single linear layer.
Training
SFT in cadrille mostly follows the training procedure of CAD-Recode. The only difference is using batch size of 8 and four gradient accumulation steps due to increase of memory for longer prompts in the multimodal scenario. The SFT model is trained with AdamW optimizer for 120k steps and learning rate of 2e-4 on a single H100 GPU.
RL fine-tuning
The RL fine-tuning hyperparameters are listed in Tab. 10 (DPO) and  11 (Dr. CPPO). In each experiment, the model is initialized with the weights of an SFT model trained on point clouds and images, and then fine-tuned only on images. All fine-tuning experiments are performed on 8 H100 GPUs.
Hyperparameter Value
Optimizer Adam
Number of epochs 20
Batch size 160
Learning rate 1e-05
KL regularization coef ($\beta$) 0.3
Table 10: DPO tuning hyperparameters.
Hyperparameter Value
Optimizer Adam
Number of epochs 20
Batch size 128
Learning rate 3e-05
Updates per batch 3
PPO $\epsilon$ 0.1
GRPO group size ($G$) 16
CPPO number of samples ($N$) 4
Table 11: RL fine-tuning hyperparameters.
Appendix D Miscellaneous
Python codes
In Fig. 10, we provide Python code with CadQuery library, produced by cadrille. When executed, these Python scripts generate CAD models. Evidently, cadrille  is not limited to basic geometric primitives from the DeepCAD dataset (such as line, arc, circle), but is also capable of producing more advanced shapes from the CAD-Recode dataset (box, rectangle, cylinder).
Refer to caption {pythoncode} import cadquery as cq w0=cq.Workplane(’XY’,origin=(0,0,11)) r=w0.sketch().segment((-100,23),(-43,-23)).segment((-63,-95)).segment((0,-55)) .segment((62,-96)).segment((42,-23)).segment((100,23)).segment((27,26)) .segment((0,96)).segment((-26,26)).close().assemble().finalize().extrude(-23)
Refer to caption {pythoncode} import cadquery as cq w0=cq.Workplane(’XY’,origin=(0,0,88)) r=w0.sketch().push([(-87,-89)]).rect(20,22).push([(-87,89)]).rect(20,22) .push([(88,-89)]).rect(20,22).push([(88,90)]).rect(20,20).finalize().extrude(-176) .union(w0.workplane(offset=-15/2).box(200,200,15))
Refer to caption {pythoncode} import cadquery as cq w0=cq.Workplane(’YZ’,origin=(-63,0,0)) r=w0.sketch().circle(33).circle(27,mode=’s’).finalize().extrude(-37) .union(w0.sketch().segment((-55,-67),(55,-67)).segment((55,-66)).arc((87,0), (55,66)).segment((-9,67)).segment((-55,67)).arc((-87,1),(-55,-67)).assemble() .push([(-53,-39)]).circle(12,mode=’s’).push([(-53,39)]).circle(11,mode=’s’) .push([(53,-43)]).circle(12,mode=’s’).finalize().extrude(21)) .union(w0.sketch().circle(34).circle(27,mode=’s’).finalize().extrude(163))
Refer to caption {pythoncode} import cadquery as cq w0=cq.Workplane(’XY’,origin=(0,0,56)) w1=cq.Workplane(’XY’,origin=(0,0,26)) r=w0.workplane(offset=-149/2).cylinder(149,18.5) .union(w0.sketch().circle(100).circle(78,mode=’s’).finalize().extrude(-111)) .union(w1.workplane(offset=-52/2).cylinder(52,77)) .union(w1.workplane(offset=6/2).moveTo(97,0).box(3.5,9.5,6))
Figure 10: cadrille predictions on DeepCAD, Fusion360, and CC3D datasets. Each row contains predicted CadQuery Python code and its result after execution in Python interpreter.
Ground truth CAD model [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Real-world noisy scan [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 11: Ground truth CAD models (top row) and noisy scans (bottom row) from the real-world CC3D dataset.
CC3D scans
Following CAD-Recode, we treat experiments with CC3D dataset as an emulation of a real-world experiment. CC3D contains scans of the physical objects paired with ground truth CAD models. As shown in Fig. 11, CC3D scans contain artifacts such as surface noise, smoothed edges, and missing parts. In our experiments, we sample points from the surfaces of these real noisy scans instead of the perfect CAD models, which essentially makes the experimental scenario much more realistic.
Besides, CC3D contains generally more complex CAD models, which are constructed using operations beyond simple extrusion, including revolution, chamfer, and fillet.